OpenStudy (anonymous):

A example of x=54x +54

OpenStudy (anonymous):

i wish i could help but im not good at math -__-.

jhonyy9 (jhonyy9):

so if you need solving for x so than x=54x +54 subtract from both sides 54 x-54=54x +54 -54 x-54=54x subtract from both sides x x-x-54=54x-x -54=53x divide both sides by 53 -54/53 =x